# On obstructions to the Euler system method for Rankin-Selberg convolutions

### v3

@@ 2 Special elements in the image of the convolution of two modular forms
[…]
    sc6special The multiplicative order of $\varepsilon_{f}$ is a power of $2$, $\varepsilon_{g}(G_{{\mathbb{Q}}})$ contains an element of odd prime order $q\mid|\tilde{\rho}_{g}(G_{{\mathbb{Q}}})|$, and, if $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})\simeq A_{4}$, then $\rho_{g}(G_{{\mathbb{Q}}})\cap\operatorname{SL}_{2}(L_{0})\rightarrow\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is onto.
@@
    $T_{p}E\otimes\rho$ has Euler-adapted Galois image.
@@ 3 Families of dihedral counter-examples: proof of Theorem Introduction
[…]
efeg=1special Lemma 3.1  Let $M,N\in\operatorname{GL}_{2}(K)$ be matrices over a field $K$ such that $\det{M}\det{N}=1$. Then $\ker\left[M\otimes N-1\right]$ is never a line.
@@
[…]
    starts3-1special If $\eta$ is a quadratic Dirichlet character, then the newform $f\otimes\eta$ has the same weight and character as $f$ and the same attached fields $H,L,F$. Hence, if $\beta$ is another quadratic Dirichlet character, the couple of newforms $(f\otimes\eta,g\otimes\beta)$ also satisfies the assumptions made at the beginning of this Section.
@@
[…]
    starts3-2special Let $\mathfrak{p}\subset\mathcal{O}_{L_{0}}$ be a prime ideal and $\sigma\in G_{{\mathbb{Q}}}$ be such that the kernel of $\sigma-1$ acting on $T_{f,g,\mathfrak{p}}$ is a line. Let $M\in\operatorname{GL}_{2}(\mathcal{O}_{(L_{0})_{\mathfrak{p}}})$ and $N\in\operatorname{GL}_{2}(\mathcal{O}_{L_{0}})$ be the matrices of the action of $\sigma$ on $T_{f,\mathfrak{p}},T_{g}$. Then one has $\varepsilon(\sigma)=\varepsilon_{f}(\sigma)\varepsilon_{g}(\sigma)=\det{M}\det{N}\neq 1$ by Lemma 3, so $\sigma\notin G_{K}$. Since $g$ has real or complex multiplication by $K$, NN has trace zero, so, if $u\in{\mathbb{C}}^{\times}$ denotes a square root of $\det{N}$, $N$ is similar to $\Delta_{u,-u}$. This implies that $N^{2}$ is a scalar matrix. Moreover, $N$ is similar to N-N, so $M\otimes N$ is similar to $M\otimes(-N)=-(M\otimes N)$, so that $\ker(M\otimes N+1)$ is also a line.
@@
[…]
    starts3-3special The discussion of (ii) has the following implication to the situation of (i): the representation $T_{f,g,\mathfrak{p}}$ satisfies (wE) if and only if $T_{f\otimes\eta,g\otimes\beta,\mathfrak{p}}$ satisfies (wE).
@@
[…]
se-equiv-Bpspecial Proposition 3.3  Let $\mathfrak{p}$ be a prime ideal of $\mathcal{O}_{L_{0}}$ of residue characteristic $p$, good for $(f,g)$. Then the following are equivalent:
@@
By Proposition 2 and Corollary 2, we can choose $\sigma_{1}\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}\cap sG_{H}$ such that $\sigma_{1}$ acts on $T_{f,\mathfrak{p}}$ by the matrix $\alpha^{-1}\varepsilon_{f}(s)\begin{pmatrix}\nu^{\prime}&1\\ 0&\nu^{\prime}\end{pmatrix}$. Since $\varepsilon(\sigma_{1})=\varepsilon(s)=-1$, the matrix $\tilde{\rho}_{g}(\sigma_{1})\in\operatorname{PGL}_{2}(\mathcal{O}_{L_{0}})$ has order two. Since $\det{\rho_{g}(\sigma_{1})}=-\varepsilon_{f}(\sigma_{1})^{-1}=-\varepsilon_{f}(s)^{-1}$ and $p\neq 2$, $\rho_{g}(\sigma_{1})$ is $\operatorname{GL}_{2}(\mathcal{O}_{(L_{0}){\mathfrak{p}}})$-conjugate to $\Delta_{\omega,-\omega}$. Therefore, $\sigma_{1}$ acts on $T_{f,g,\mathfrak{p}}$ by $\begin{pmatrix}1&a&0&0\\ 0&1&0&0\\ 0&0&-1&\ast\\ 0&0&0&-1\end{pmatrix}$ with $a\in\mathcal{O}_{(L_{0})_{\mathfrak{p}}}^{\times}$: hence $T_{f,g,\mathfrak{p}}/(\sigma_{1}-1)T_{f,g,\mathfrak{p}}$ is free of $\mathcal{O}_{(L_{0})_{\mathfrak{p}}}$-rank one. $\Box$
[…]
bidualspecial Corollary 3.4  The answer to Question Introduction for $(f,g)$ is negative if, and only if, $\varepsilon(M)=1$.
Proof. –  By Chebotarev and the previous Proposition, the answer to Question Introduction for $(f,g)$ is positive if and only if $\varepsilon$ cannot be written as a $\mathcal{B}(\cdot,\tau)$ for any $\tau\in G_{F}$. Because $\mathcal{B}$ is bilinear with values in a ${\mathbb{F}}_{2}$-vector space, the set $\{\mathcal{B}(\cdot,\tau)\mid\tau\in G_{F}\}$ is exactly the set of homomorphisms $f:\mathrm{Gal}(H/{\mathbb{Q}})\rightarrow\{\pm 1\}$ such that $f(M)=1$, and the conclusion follows. $\Box$
@@
[…]
whenepsilonexistsspecial Proposition 3.5  Let $H_{0}\subset H$ be the subfield such that $G_{H_{0}}=\ker{\varepsilon_{f}}$. Let $c\in\mathrm{Gal}(H/{\mathbb{Q}})$ denote the complex conjugation. Assume that one of the following holds:
@@ 1. $(a)$
[…]
    weps-aspecial $k$ is even and $H$ is not totally real,
@@ 3. $(c)$
[…]
    weps-cspecial $k$ is odd, $H:H_{0}]=2$, and $\varepsilon_{f}$ does not take its values in $L^{\times 2}$ (this last condition is satisfied if $L$ does not contain a fourth root of unity $i$),
@@ 4. $(d)$
[…]
    weps-dspecial $k$ is odd, $H:H_{0}]=2$, $i\in L$, and for all $(\gamma,\delta)\in\Gamma_{f}$, $\gamma(i)=\delta(-1)i$,
@@ 5. $(e)$
[…]
    weps-especial $k$ is odd, $\mathrm{Gal}(H/H_{0})$ contains a unique non-trivial element $\sigma_{0}$, $i\in L$ and, for some $(\gamma,\delta)\in\Gamma_{f}$, one has $\gamma(i)\neq\delta(\sigma_{0})\delta(-1)i$.
@@
Using Corollary 3, this Proposition implies:
Corollary 3.6  Suppose that the hypotheses of Proposition 3 hold: by Proposition 3, there exists a non-trivial character $\varepsilon:\mathrm{Gal}(H/{\mathbb{Q}})/M\rightarrow\{\pm 1\}$ mapping the complex conjugation to $(-1)^{k-1}$. Let $g$ be a weight one newform with character $\varepsilon\varepsilon_{f}^{-1}$ which has either complex multiplication by $\varepsilon$ if $k$ is even, or real multiplication by $\varepsilon$ is $k$ is odd. Then there are infinitely many prime ideals $\mathfrak{p}\subset L_{0}$ (with residue characteristic $p$) such that there is no $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$ such that the kernel of the action of $\sigma-1$ on $V_{f,g,\mathfrak{p}}$ is a line.
@@
    For a newform $f\in\mathcal{S}_{3}(\Gamma_{1}(24))$ with LMFDB label $24.3.h.c$, the condition $(c)$ of Proposition 3 is satisfied: indeed, in this situation $L={\mathbb{Q}}(\sqrt{2},\sqrt{-7})$ does not contain $i$, while $H={\mathbb{Q}}(\sqrt{2},\sqrt{-3})$, and $H_{0}={\mathbb{Q}}(\sqrt{-6})$ is such that $H:H_{0}]=2$. Indeed, $B$ contains the two distinct non-trivial elements $\mathcal{B}(c,\cdot)$ and $\mathcal{B}(\mathrm{Frob}_{5},\cdot)$ (they are the Kummer characters associated to $\sqrt{14}$ and $\sqrt{2}$ respectively), so that $M$ is trivial and $\mathrm{Gal}(H/{\mathbb{Q}})$ is larger than $\{\mathrm{id},c\}$.
@@
When $K$ is imaginary, we may assume (after multiplying $z$ with a rational number) that $z=(z_{0},z_{f})$ with $z_{f}\in\hat{{\mathbb{Z}}}^{\times}$ and $z_{0}\in{\mathbb{R}}^{\times}_{+}$, and thus that its image in $\mathbb{A}_{K}^{\times}$ is contained in $\mathcal{O}_{K}^{\times}\prod_{v}{M_{v}}$. It is then enough to find $D$ such that any unit $u\in\mathcal{O}_{K}^{\times}$ which is congruent to an integer modulo $D$ is congruent to $\pm 1$ modulo $C$ (this $\pm 1$ corresponds to the fact that the complex conjugation is contained in the kernel). Since $\mathcal{O}_{K}^{\times}$ has order $2,4$ or $6$, $D=11C$ works, since ${\mathbb{F}}_{11}$ does not contain primitive third or fourth roots of unity.
@@ 4 Examples and counter-examples
[…]
sect-counterexspecial So far, we have identified in Theorem Introduction and in Section 2 some sufficient conditions to ensure a positive answer to Question Introduction. On the other hand, the situation studied in Section 3 provides in some cases families of counter-examples. The purpose of this Section is to study some situations where the results of neither Section apply. We provide a few counter-examples which do not stem from Section 3. We also showcase additional examples proving that, while the exceptional cases highlighted in Propositions 2 and 2 can occur, they are not necessarily incompatible with Euler-adapted Galois images. Some of these counter-examples are explicit: as in the examples of Section 3, we will then always refer to newforms by their LMFDB label [14]. While this label only refers to the equivalence class of the modular forms under the action of $G_{{\mathbb{Q}}}$, the arguments work regardless of the choice of embedding of the modular forms in $\mathbb{C}$}.
[…]
special-counterspecial Proposition 4.1  Assume that $\mathfrak{p}$ is a good prime, and that $(f,g,p)$ falls in one of the following cases:
@@
The proof strategy is the same in all cases: using the LMFDB data, we check that in every case, $F={\mathbb{Q}}$, $L,H$ are quadratic number fields, $f$ has trivial character, and $\mathfrak{p}$ is coprime to $|\rho_{g}(G_{{\mathbb{Q}}})|$. Then, for each coset $C=\sigma G_{H(\mu_{p^{\infty}})}\subset G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$, we can determine an element $\alpha\in\mathcal{O}_{L,\mathfrak{p}}^{\times}$ attached to $C$ by Corollary 2, since we know that $\alpha^{2}\in F$.
@@
Assuming that $\varepsilon_{f}$ is the trivial character, that $F={\mathbb{Q}}$, and that $\mathfrak{p}$ is coprime to $2|\rho_{g}(G_{{\mathbb{Q}}})|$, we have thus proved the following. Pick, for every coset $C$ of $G_{H(\mu_{p^{\infty}})}$ in $G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$, some $\alpha_{C}\in\mathcal{O}_{L,\mathfrak{p}}^{\times}$ given by Corollary 2, and let $V_{C}\subset L_{0}$ be a set containing the eigenvalues of non-scalar matrices $N\in\rho_{g}(C)$ such that $\det{N}\neq 1$. If, for each $C$ such that $\alpha_{C}\notin{\mathbb{Z}}_{p}^{\times}$, $V_{C}$ does not contain an element $u$ such that $1\in\{u^{p-1},u^{p+1}\}$, and, for each $C$ such that $\alpha_{C}\in{\mathbb{Z}}_{p}^{\times}$, $V_{C}$ does not contain an element such that $1\in\{u^{p-1},u^{p+1}\}$, then $T_{f,g,\mathfrak{p}}$ does not satisfy (wE).
@@
The LMFDB data stipulates that $\ker{\tilde{\rho}_{g}}\not\subset G_{H}$, so $\tilde{\rho}_{g}(G_{H})=\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$. Using MAGMA [1], we can also see that $\ker{\rho_{g}}\subset G_{H}$, so that $\rho_{g}(G_{{\mathbb{Q}}}):\rho_{g}(G_{H})]=2$.
For the sake of easier notation, write $C\subset\rho_{g}(G_{{\mathbb{Q}}})$ for the pre-image of the non-trivial elements of the cyclic subgroup of order $3$ of $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$.
Since $\rho_{g}(G_{H})$ and $\rho_{g}(G_{{\mathbb{Q}}})$ have the same determinant and projective image, we can split elements of $\rho_{g}(G_{{\mathbb{Q}}})$ in six categories: $\{\pm iI_{2},\pm I_{2}\}$, scalar matrices of order $8$, matrices in $C$ with determinant $\pm 1$, matrices in $C$ with determinant $\pm i$, matrices outside $C$ with determinant $\pm 1$, matrices outside $C$ with determinant $\pm i$. Then $\rho_{g}(G_{H})$ is exactly made with the matrices of the first, third, and last categories.
@@
Let $N\in C$ be a matrix in the fourth category. Then $\det{N}=\pm i$ and $N^{3}$ is scalar. Then $N\sim\omega\Delta_{j,j^{2}}$ for some primitive $8$-th root of unity $\omega$, so any eigenvalue $u$ of $N$ has order $24$. Since $p\not\equiv\pm 1\pmod{12}$, none of $u^{2(p\pm 1)}\neq 1$ so $1\notin\{u^{p\pm 1}\}$, and we are done.
Let $N$ be a matrix in the fifth category such that $\det{N}\neq 1$. Then $\det{N}=-1$ and the projective image of $N$ has order $2$, so the eigenvalues of $N$ are $\pm 1$, $1^{p\pm 1}=(-1)^{p\pm 1}=1$, so we are done.
@@
[…]
pvaries-finespecial Proposition 4.2  Let $f\in\mathcal{S}_{k}(\Gamma_{1}(N)),g\in\mathcal{S}_{1}(\Gamma_{1}(M))$ be newforms with $k\geq 2$ and respective characters $\varepsilon_{f},\varepsilon_{g}$. Assume that $f$ has no complex multiplication and let $H,L,F$ be the fields attached to $f$ as in Section 1. Let $L_{0}\subset{\mathbb{C}}$ be a number field containing $L$, the Fourier coefficients of $g$ and the eigenvalues of all matrices in $\rho_{g}(G_{{\mathbb{Q}}})$. Assume furthermore that:
@@
By the second part of Proposition 2, we can fix an element $\tau\in G_{H_{0}(\mu_{p^{\infty}})}$ such that $\rho_{g}(\tau)\notin\rho_{g}(G_{H})$. Let $\alpha\in L^{\times}\cap\mathcal{O}_{L,\mathfrak{p}}^{\times}$ be the element attached to $\tau$ by Corollary 2. Then $\alpha^{2}\in\mathcal{O}_{F,\mathfrak{p}}^{\times}$ and $\alpha\notin F$. Let $\mathfrak{p}\subset\mathcal{O}_{L_{0}}$ be a good prime for $(f,g)$ of residue characteristic $p$. We will show that the representation $T_{f,g,\mathfrak{p}}$ satisfies (sE) by distinguishing between the following three cases:
@@
[…]
    pvaries-1special if $2$ is a square in $F_{\mathfrak{p}}$,
@@
[…]
    pvaries-2special if $\alpha^{2}\in F$ is a square in $F_{\mathfrak{p}}$,
@@
[…]
    pvaries-3special if $2\alpha^{2}$ is a square in $F_{\mathfrak{p}}$.
@@
[…]
special2-exc-notholdspecial Corollary 4.3  Let $f\in\mathcal{S}_{2}(\Gamma_{0}(63)),g\in\mathcal{S}_{1}(\Gamma_{1}(675))$ be newforms with respective labels $63.2.a.b$ and $675.1.g.a$. Then $T_{f,g,\mathfrak{p}}$ has Euler-adapted Galois image for every good prime $\mathfrak{p}$.

### v4

@@ 2 Special elements in the image of the convolution of two modular forms
    The multiplicative order of $\varepsilon_{f}$ is a power of $2$, $\varepsilon_{g}(G_{{\mathbb{Q}}})$ contains an element of odd prime order $q\mid|\tilde{\rho}_{g}(G_{{\mathbb{Q}}})|$, and, if $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})\simeq A_{4}$, then $\rho_{g}(G_{{\mathbb{Q}}})\cap\operatorname{SL}_{2}(L_{0})\rightarrow\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is onto.
@@
    $T_{p}E\otimes\rho\simeq T_{f,g,\mathfrak{p}}$ has Euler-adapted Galois image.
@@ 3 Families of dihedral counter-examples: proof of Theorem Introduction
Lemma 3.1  Let $M,N\in\operatorname{GL}_{2}(K)$ be matrices over a field $K$ such that $\det{M}\det{N}=1$. Then $\ker\left[M\otimes N-1\right]$ is never a line.
@@
    If $\eta$ is a quadratic Dirichlet character, then the newform $f\otimes\eta$ has the same weight and character as $f$ and the same attached fields $H,L,F$. Hence, if $\beta$ is another quadratic Dirichlet character, the couple of newforms $(f\otimes\eta,g\otimes\beta)$ also satisfies the assumptions made at the beginning of this Section.
@@
    Let $\mathfrak{p}\subset\mathcal{O}_{L_{0}}$ be a prime ideal and $\sigma\in G_{{\mathbb{Q}}}$ be such that the kernel of $\sigma-1$ acting on $T_{f,g,\mathfrak{p}}$ is a line. Let $M\in\operatorname{GL}_{2}(\mathcal{O}_{(L_{0})_{\mathfrak{p}}})$ and $N\in\operatorname{GL}_{2}(\mathcal{O}_{L_{0}})$ be the matrices of the action of $\sigma$ on $T_{f,\mathfrak{p}},T_{g}$. Then one has $\varepsilon(\sigma)=\varepsilon_{f}(\sigma)\varepsilon_{g}(\sigma)=\det{M}\det{N}\neq 1$ by Lemma 3, so $\sigma\notin G_{K}$. Since $g$ has real or complex multiplication by $K$, NN has trace zero, so, if $u\in{\mathbb{C}}^{\times}$ denotes a square root of $\det{N}$, $N$ is similar to $\Delta_{u,-u}$. This implies that $N^{2}$ is a scalar matrix. Moreover, $N$ is similar to N-N, so $M\otimes N$ is similar to $M\otimes(-N)=-(M\otimes N)$, so that $\ker(M\otimes N+1)$ is also a line.
@@
    The discussion of (ii) has the following implication to the situation of (i): the representation $T_{f,g,\mathfrak{p}}$ satisfies (wE) if and only if $T_{f\otimes\eta,g\otimes\beta,\mathfrak{p}}$ satisfies (wE).
@@
Proposition 3.3  Let $\mathfrak{p}$ be a prime ideal of $\mathcal{O}_{L_{0}}$ of residue characteristic $p$, good for $(f,g)$. Then the following are equivalent:
@@
By Proposition 2 and Corollary 2, we can choose $\sigma_{1}\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}\cap sG_{H}$ acting on $T_{f,\mathfrak{p}}$ by the matrix $\alpha^{-1}\varepsilon_{f}(s)\begin{pmatrix}\nu^{\prime}&1\\ 0&\nu^{\prime}\end{pmatrix}$. Since $\varepsilon(\sigma_{1})=\varepsilon(s)=-1$, the matrix $\tilde{\rho}_{g}(\sigma_{1})\in\operatorname{PGL}_{2}(\mathcal{O}_{L_{0}})$ has order two. Since $\det{\rho_{g}(\sigma_{1})}=-\varepsilon_{f}(\sigma_{1})^{-1}=-\varepsilon_{f}(s)^{-1}$ and $p\neq 2$, $\rho_{g}(\sigma_{1})$ is $\operatorname{GL}_{2}(\mathcal{O}_{(L_{0}){\mathfrak{p}}})$-conjugate to $\Delta_{\omega,-\omega}$. Therefore, $\sigma_{1}$ acts on $T_{f,g,\mathfrak{p}}$ by $\begin{pmatrix}1&a&0&0\\ 0&1&0&0\\ 0&0&-1&\ast\\ 0&0&0&-1\end{pmatrix}$ with $a\in\mathcal{O}_{(L_{0})_{\mathfrak{p}}}^{\times}$: hence $T_{f,g,\mathfrak{p}}/(\sigma_{1}-1)T_{f,g,\mathfrak{p}}$ is free of $\mathcal{O}_{(L_{0})_{\mathfrak{p}}}$-rank one. $\Box$
Corollary 3.4  The answer to Question Introduction for $(f,g)$ is negative if, and only if, $\varepsilon(M)=1$.
Proof. –  By Chebotarev and Proposition 3, the answer to Question Introduction for $(f,g)$ is positive if and only if $\varepsilon$ cannot be written as a $\mathcal{B}(\cdot,\tau)$ for any $\tau\in G_{F}$. Because $\mathcal{B}$ is bilinear with values in a ${\mathbb{F}}_{2}$-vector space, the set $\{\mathcal{B}(\cdot,\tau)\mid\tau\in G_{F}\}$ is exactly the set of homomorphisms $f:\mathrm{Gal}(H/{\mathbb{Q}})\rightarrow\{\pm 1\}$ such that $f(M)=1$, and the conclusion follows. $\Box$
@@
Proposition 3.5  Let $H_{0}\subset H$ be the subfield such that $G_{H_{0}}=\ker{\varepsilon_{f}}$. Let $c\in\mathrm{Gal}(H/{\mathbb{Q}})$ denote the complex conjugation. Assume that one of the following holds:
@@ 1. $(a)$
    $k$ is even and $H$ is not totally real,
@@ 3. $(c)$
    $k$ is odd, $H:H_{0}]=2$, and $\varepsilon_{f}$ does not take its values in $L^{\times 2}$ (this last condition is satisfied if $L$ does not contain a fourth root of unity $i$),
@@ 4. $(d)$
    $k$ is odd, $H:H_{0}]=2$, $i\in L$, and for all $(\gamma,\delta)\in\Gamma_{f}$, $\gamma(i)=\delta(-1)i$,
@@ 5. $(e)$
    $k$ is odd, $\mathrm{Gal}(H/H_{0})$ contains a unique non-trivial element $\sigma_{0}$, $i\in L$ and, for some $(\gamma,\delta)\in\Gamma_{f}$, one has $\gamma(i)\neq\delta(\sigma_{0})\delta(-1)i$.
@@
By Corollary 3, Proposition 3 implies:
Corollary 3.6  Suppose that the hypotheses of Proposition 3 hold: by Proposition 3, there exists a non-trivial character $\varepsilon:\mathrm{Gal}(H/{\mathbb{Q}})/M\rightarrow\{\pm 1\}$ mapping the complex conjugation to $(-1)^{k-1}$. Let $g$ be a weight one newform with character $\varepsilon\varepsilon_{f}^{-1}$ which has either complex multiplication by $\varepsilon$ if $k$ is even, or real multiplication by $\varepsilon$ is $k$ is odd. Then there are infinitely many maximal ideals $\mathfrak{p}\subset\mathcal{O}_{L_{0}}$ (with residue characteristic $p$) such that there is no $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$ such that the kernel of the action of $\sigma-1$ on $V_{f,g,\mathfrak{p}}$ is a line.
@@
    For a newform $f\in\mathcal{S}_{3}(\Gamma_{1}(24))$ with LMFDB label $24.3.h.c$, the condition $(c)$ of Proposition 3 is satisfied: indeed, in this situation $L={\mathbb{Q}}(\sqrt{2},\sqrt{-7})$ does not contain $i$, while $H={\mathbb{Q}}(\sqrt{2},\sqrt{-3})$ and $H_{0}={\mathbb{Q}}(\sqrt{-6})$ satisfy $H:H_{0}]=2$. Indeed, $B$ contains the two distinct non-trivial elements $\mathcal{B}(c,\cdot)$ and $\mathcal{B}(\mathrm{Frob}_{5},\cdot)$ (they are the Kummer characters associated to $\sqrt{14}$ and $\sqrt{2}$ respectively), so $M$ is trivial and $\mathrm{Gal}(H/{\mathbb{Q}})$ properly contains $\{\mathrm{id},c\}$.
@@
When $K$ is imaginary, we may assume (after multiplying $z$ with a rational number) that $z=(z_{0},z_{f})$ with $z_{f}\in\hat{{\mathbb{Z}}}^{\times}$ and $z_{0}\in{\mathbb{R}}^{\times}_{+}$, and thus that its image in $\mathbb{A}_{K}^{\times}$ is contained in $\mathcal{O}_{K}^{\times}\prod_{v}{M_{v}}$. It is then enough to find $D$ such that any unit $u\in\mathcal{O}_{K}^{\times}$ which is congruent to an integer modulo $D$ is congruent to $\pm 1$ modulo $C$ (this $\pm 1$ corresponds to the fact that the complex conjugation is contained in the kernel). Since $\mathcal{O}_{K}^{\times}$ has order $2,4$ or $6$, $D=11C$ works, since ${\mathbb{F}}_{11}$ does not contain any primitive third or fourth root of unity.
@@ 4 Examples and counter-examples
So far, we have identified in Theorem Introduction and in Section 2 some sufficient conditions to ensure a positive answer to Question Introduction. On the other hand, the situation studied in Section 3 provides in some cases families of counter-examples. The purpose of this section is to study some situations where these results do not apply. We provide a few counter-examples which do not stem from Section 3. We also showcase additional examples proving that, while the exceptional cases highlighted in Propositions 2 and 2 can occur, they are not necessarily incompatible with Euler-adapted Galois images. Some of these counter-examples are explicit: as in the examples of Section 3, we will then always refer to newforms by their LMFDB label [14]. While this label only refers to the equivalence class of the modular forms under the action of $G_{{\mathbb{Q}}}$, the arguments work regardless of the choice of embedding of the modular forms in $\mathbb{C}$}.
Proposition 4.1  Assume that $\mathfrak{p}$ is a good prime, and that $(f,g,p)$ falls in one of the following cases:
@@
The proof strategy is the same in all cases: using the LMFDB data, we check that in every case, $F={\mathbb{Q}}$, $L,H$ are quadratic number fields, $f$ has trivial character, and $\mathfrak{p}$ is good and coprime to $|\rho_{g}(G_{{\mathbb{Q}}})|$. Then, for each coset $C=\sigma G_{H(\mu_{p^{\infty}})}\subset G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$, we can determine an element $\alpha\in\mathcal{O}_{L,\mathfrak{p}}^{\times}$ attached to $C$ by Corollary 2, since we know that $\alpha^{2}\in F$.
@@
Assuming that $\varepsilon_{f}$ is the trivial character, that $F={\mathbb{Q}}$, and that $\mathfrak{p}$ is coprime to $2|\rho_{g}(G_{{\mathbb{Q}}})|$, we have thus proved the following. Pick, for every coset $C$ of $G_{H(\mu_{p^{\infty}})}$ in $G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$, some $\alpha_{C}\in\mathcal{O}_{L,\mathfrak{p}}^{\times}$ given by Corollary 2, and let $V_{C}\subset L_{0}$ be a set containing the eigenvalues of non-scalar matrices $N\in\rho_{g}(C)$ such that $\det{N}\neq 1$. If, for each $C$ such that $\alpha_{C}\notin{\mathbb{Z}}_{p}^{\times}$, $V_{C}$ does not contain an element $u$ such that $1\in\{u^{p-1},u^{p+1}\}$, and, for each $C$ such that $\alpha_{C}\in{\mathbb{Z}}_{p}^{\times}$, $V_{C}$ does not contain an element $u$ such that $1\in\{u^{p-1},u^{p+1}\}$, then $T_{f,g,\mathfrak{p}}$ does not satisfy (wE).
@@
By the LMFDB data, one has $G_{H}\subsetneq G_{H}\cdot\ker{\tilde{\rho}_{g}}$, so $\tilde{\rho}_{g}(G_{H})=\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$. Using MAGMA [1], we can also see that $\ker{\rho_{g}}\subset G_{H}$, so that $\rho_{g}(G_{{\mathbb{Q}}}):\rho_{g}(G_{H})]=2$.
For the sake of easier notation, write $C\subset\rho_{g}(G_{{\mathbb{Q}}})$ for the pre-image of the cyclic subgroup of order $3$ of $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$.
Since $\rho_{g}(G_{H})$ and $\rho_{g}(G_{{\mathbb{Q}}})$ have the same determinant and projective image, we can split elements of $\rho_{g}(G_{{\mathbb{Q}}})$ in six categories: $\{\pm iI_{2},\pm I_{2}\}$, scalar matrices of order $8$, non-scalar matrices in $C$ with determinant $\pm 1$, non-scalar matrices in $C$ with determinant $\pm i$, matrices outside $C$ with determinant $\pm 1$, matrices outside $C$ with determinant $\pm i$. Then $\rho_{g}(G_{H})$ is exactly made with the matrices of the first, third, and last categories.
@@
Let $N\in C$ be a matrix in the fourth category. Then $\det{N}=\pm i$ and $N^{3}$ is scalar. Then $N\sim\omega\Delta_{j,j^{2}}$ for some primitive $8$-th root of unity $\omega$, so any eigenvalue $u$ of $N$ has order $24$. Since $p\not\equiv\pm 1\pmod{12}$, one has $1\notin\{u^{2(p\pm 1)}\}$, so $1\notin\{u^{p\pm 1}\}$, and we are done.
Let $N$ be a matrix in the fifth category such that $\det{N}\neq 1$. Then $\det{N}=-1$ and the projective image of $N$ has order $2$, and we conclude as in (a).
@@
Proposition 4.2  Let $f\in\mathcal{S}_{k}(\Gamma_{1}(N)),g\in\mathcal{S}_{1}(\Gamma_{1}(M))$ be newforms with $k\geq 2$ and respective characters $\varepsilon_{f},\varepsilon_{g}$. Assume that $f$ has no complex multiplication and let $H,L,F$ be the fields attached to $f$ as in Section 1. Let $L_{0}\subset{\mathbb{C}}$ be a number field containing $L$, the Fourier coefficients of $g$ and the eigenvalues of all matrices in $\rho_{g}(G_{{\mathbb{Q}}})$. Assume furthermore that:
@@
By the second part of Proposition 2, we can fix an element $\tau\in G_{H_{0}(\mu_{p^{\infty}})}$ such that $\rho_{g}(\tau)\notin\rho_{g}(G_{H})$. Let $\mathfrak{p}\subset\mathcal{O}_{L_{0}}$ be a good prime for $(f,g)$ of residue characteristic $p$. Let $\alpha\in\mathcal{O}_{L,\mathfrak{p}}^{\times}$ be the element attached to $\tau$ by Corollary 2. Then $\alpha^{2}\in\mathcal{O}_{F,\mathfrak{p}}^{\times}$ and $\alpha\notin F$. We will show that the representation $T_{f,g,\mathfrak{p}}$ satisfies (sE) by distinguishing between the following three cases:
@@
    if $2$ is a square in $F_{\mathfrak{p}}$,
@@
    if $\alpha^{2}\in F$ is a square in $F_{\mathfrak{p}}$,
@@
    if $2\alpha^{2}$ is a square in $F_{\mathfrak{p}}$.
@@
Corollary 4.3  Let $f\in\mathcal{S}_{2}(\Gamma_{0}(63)),g\in\mathcal{S}_{1}(\Gamma_{1}(675))$ be newforms with respective labels $63.2.a.b$ and $675.1.g.a$. Then $T_{f,g,\mathfrak{p}}$ has Euler-adapted Galois image for every good prime $\mathfrak{p}$.